### v1

@@ 1 Introduction
More broadly, our work takes a step towards assessing the utility of both learning agents in decentralized, misaligned environments. Our model and results capture the general setting where the player utilities to be arbitrarily related, where players might not even agree upon which pairs of actions give similar or different rewards. This motivated us to design benchmarks which are tolerant to small errors in the other player. We hope that our benchmarks and algorithms serve as a starting point for assessing when two-agent learning systems in misaligned environments can ensure high utility for both agents.
@@ 1.1 Related Work
Several papers study decentralized online learning in SGs. Camara et al. (2020); Collina et al. (2023b) posit that the follower runs a no-counterfactual-internal-regret algorithm and design no-regret algorithms for the leader. However, they assume strong alignment between the players’ rewards Camara et al. (2020) requires that a follower choosing an $\epsilon$-suboptimal action only results in an $O(\epsilon)$ utility loss for the leader. [^7] See Assumption 2 in Camara et al. (2020). Appendix D therein considers some relaxations, but they lead to $\Omega(T)$ worst-case regret. Collina et al. (2023b) partially relax this assumption, but still require the existence of stable actions for the leader. In contrast, we do not place any alignment conditions: in fact, misalignement is the driver of our linear regret result for the original Stackelberg benchmarks (Theorem 3.1). Other differences are that we focus on stochastic, rather than adversarial, rewards, and our benchmark is independent of the follower’s choice of learning algorithm. [^8] However, our regret bounds assume that the follower’s algorithm gracefully improves over time, see Section 2.3. Haghtalab et al. (2023) take a different perspective and consider the follower running a calibrated algorithm. They design a leader algorithm which waits for the follower to partially converge, and show that the Stackelberg value is obtained in the limit as $T\rightarrow\infty$. In contrast, we focus on instance-independent regret bounds for a fixed time horizon $T$, which requires us to relax the benchmark. Other differences are we focus on stochastic, rather than deterministic, rewards, we assume the follower observes the leader’s action, and we consider the follower’s utility in addition to the leader’s utility.
@@
The literature on learning in SGs is vast and includes many other variations. Many works (e.g., (Letchford et al., 2009; Balcan et al., 2015; Zhao et al., 2023)) consider the leader performing (offline or) online learning and followers myopically best-responding. Other model variants studied include the leader strategizing against a follower who is running a no-regret learning algorithm (Braverman et al., 2018; Deng et al., 2019; Guruganesh et al., 2024; Brown et al., 2023), the leader and follower both running gradient-based algorithms (Fiez et al., 2019), non-myopic followers who best-respond to a discounted utility over future time steps (Haghtalab et al., 2022; Hajiaghayi et al., 2023), repeated game formulations under complete information (Zuo and Tang, 2015; Collina et al., 2023a) the leader offering a menu of actions to the follower (Han et al., 2023), and both players having side information (Harris et al., 2024). Other works have studied learning in structured SGs, including delegated choice (e.g., (Kleinberg and Kleinberg, 2018; Hajiaghayi et al., 2023)), strategic classification (e.g., (Dong et al., 2018; Chen et al., 2020; Zrnic et al., 2021; Ahmadi et al., 2021)), pricing under buyer and seller uncertainty (e.g., (Guo et al., 2023)), contract theory (e.g., (Zhu et al., 2023)), and aligned utilities (e.g., (Kao et al., 2022)).
@@
Our work also connects to a broader literature on interacting learners. This literature examines interactions between multiple bandit learners, studying aspects such as the convergence of systems of no-regret learners to coarse correlated and correlated equilibrium (e.g. (Daskalakis et al., 2011, 2021; Anagnostides et al., 2022)), multiple bandit learners competing for market share (e.g., (Aridor et al., 2020; Jagadeesan et al., 2023)), and multiple autobidding algorithms competing in an auction (e.g., (Borgs et al., 2007; Balseiro and Gur, 2019; Lucier et al., 2023)). Most closely related to this paper is corralling bandit algorithms (e.g., (Agarwal et al., 2017; Pacchiano et al., 2020)), where a “master algorithm” dynamically chooses among several “base algorithms”: our decentralized learning environment in the case of aligned player utilities is essentially an instance of corralling bandits, with the “base algorithms” corresponding to different leader actions. The interacting learner literature also examines human-algorithm collaboration studying aspects such as misalignment between engagement metrics and user welfare (e.g., (Ekstrand and Willemsen, 2016; Milli et al., 2021; Stray et al., 2021; Kleinberg et al., 2022)), impact of underspecification on human-AI misalignment (e.g., (Zhuang and Hadfield-Menell, 2020)), and “assistive” algorithmic tools (e.g. (Chan et al., 2019)). Most closely related to this paper is work on online learning in subset selection and conformal prediction Straitouri and Rodriguez (2023); Corvelo Benz and Rodriguez (2024); Straitouri et al. (2023); Wang et al. (2022); Donahue et al. (2024); Agarwal and Brown (2023), often with the goal of achieving complementarity (Bansal et al., 2021). The related area of human-AI interaction (see (Preece et al., 1994; Kim, 2015; MacKenzie, 2024; Lazar et al., 2017) for textbook treatments) studies similar questions, often from a more behavioral angle. More broadly, the interacting learner literature also studies applied domains including multi-agent reinforcement learning (see Zhang et al. (2019) for a survey) and federated learning (see Yang et al. (2019) for a survey).
@@ 2 Model and assumptions
In this section we describe our formal model. We first define an instance $\mathcal{I}=(\mathcal{A},\mathcal{B},v_{1},v_{2})$ in our setup, which captures the setup of the underlying static Stackelberg game. Let $\mathcal{A}$ be the action set for the leader (Player 1) and let $\mathcal{B}$ be the action set for the follower (Player 2). Let $v_{i}(a,b)\in[0,1]$ denote Player $i$’s value (i.e., mean reward) for the leader choosing $a$ and the follower choosing $b$. The Stackelberg equilibrium takes the following form. Let $b^{*}(a)$ be the best-response with respect to the follower’s rewards: [^9] Different variants (e.g. Strong Stackelberg Equilibrium and Weak Stackelberg Equilibrium) handle exact ties in slightly different ways. We take $b^{*}(a)$ to be the best-response with lowest utility for the leader. $b^{*}(a)=\text{argmax}_{b\in\mathcal{B}}v_{2}(a,b)$. The Stackelberg equilibrium $(a^{*},b^{*})$ is defined to be the best action the leader can take, assuming that the follower will exactly best-respond:
@@
In this paper, we move from the static Stackelberg Equilibrium environment to a repeated dynamic environment over $T$ time steps where each player selects actions using a multi-armed bandit algorithm. The environment is sequential: at each time step $t$, the leader chooses an action $a_{t}\in\mathcal{A}$ and then the follower chooses an action $b_{t}\in\mathcal{B}$. This environment is also decentralized: each player $i$ can observe their own stochastic rewards but not the stochastic rewards of the other player. We measure regret for each player $i$ by their cumulative reward across all time steps relative to a benchmark.
@@ 2.1 Interaction between players
The interaction between the leader and follower proceeds as follows. The leader chooses an algorithm $\texttt{ALG}_{1}$ mapping their history (formalized below) of observed actions and rewards to a distributions over actions $\mathcal{A}$, and the follower similarly chooses an algorithm $\texttt{ALG}_{2}$ mapping the leader’s action and their history to a distribution over actions $\mathcal{B}$. After the players select algorithms, the interaction between the leader and the follower proceeds as follows at each time step $t$:
@@
This interaction captures that the players are dynamic in their learning: in particular, this framework is sufficiently general to capture a wide range of learning strategies. However, we do not study the meta-game where agents strategically pick learning algorithms. [^11] An example of such a meta-game is Kolumbus and Nisan (2022), where each player maximizes their worst-case reward against a class of learning algorithms that the other player could choose. We believe that this model captures many real-world environments such as user-chatbot interactions and recommender system-user interactions, where agents learn about their incurred rewards from past interactions even if they do not actively optimize the higher-order manner in which they learn. See Section 2.4 (Example 2.1 and Example 2.2) for more details of how these real-world examples are captured by our model.
@@
Having described how the players interact, we next discuss the players’ history, which further enforces decentralization. Each player $i$ can only observe their own reward $r_{i,t}(a_{t},b_{t})$ (and cannot observe the reward of the other player). In the strongly decentralized setting, the follower can observe the leader’s action $a_{t}$, but the leader cannot observe the follower’s action $b_{t}$, whereas in the weakly decentralized setting, the leader can additionally observe the follower’s action $b_{t}$. One motivation for the strongly decentralized setting is interpretability: the leader and follower may be taking actions in spaces that are not mutually understandable (e.g. a chatbot’s representation of human preferences may not be interpretable). Notation for each player’s histories are presented in Appendix B.
@@ 2.2 Measuring regret
Our goal is to obtain sublinear worst-case regret for both players: that is, we will assess $\max(R_{1}(T),R_{2}(T))$. We note that this challenging goal is a departure from previous work which has typically focused solely on sublinear regret for the leader (see Section 1.1). Our motivation for selecting this objective is that (a) a human could be either the leader or the follower, and (b) the societal welfare may demand that we care about the utility of both the leader and the follower (discussed further in Section 2.4).
@@ 2.3 Assumptions on the follower’s algorithm $\texttt{ALG}_{2}$
When we analyze regret in Sections 4-5, many of our constructions do not require that the follower run a particular algorithm, but instead allow the follower to run any algorithm that has sufficiently good performance along certain fine-grained performance metrics that capture the extent to which an algorithm’s performance gracefully improves over time.
@@
In our setting, we will require similar properties to hold for the follower’s algorithm $\texttt{ALG}_{2}$, but we take account how the algorithm $\texttt{ALG}_{2}$ depends on the action $a_{t}$ which is selected by the leader’s algorithm $\texttt{ALG}_{1}$. Let $n_{t+1}(a)$ be the number of times that arm $a$ has been pulled up prior to the $(t+1)$th time step. An algorithm $\texttt{ALG}_{2}$ satisfies a high-probability instantaneous regret bound of $g$ if for any $\texttt{ALG}_{1}$ chosen by the leader and any $\mathcal{I}=(\mathcal{A},\mathcal{B},v_{1},v_{2})$, it holds that:
$\mathbb{P}\left[\forall t\in[T],a\in\mathcal{A}\mid v_{2}(a_{t},b_{t})\geq\max_{b\in\mathcal{B}}v_{2}(a_{t},b)-g(n_{t+1}(a)+1,T,\mathcal{B})\right]\geq 1-|\mathcal{A}|\cdot T^{-3},$
@@
$\mathbb{P}\left[\forall t\in[T],a\in\mathcal{A}\mid\sum_{t^{\prime}\leq t\mid a_{t^{\prime}}=a}v_{2}(a,b_{t^{\prime}})\geq\sum_{t^{\prime}\leq t\mid a_{t^{\prime}}=a}\max_{b\in\mathcal{B}}v_{2}(a,b)-h(n_{t+1}(a),T,\mathcal{B})\right]\geq 1-|\mathcal{A}|\cdot T^{-3},$
We discuss these performance metrics in more detail (including their connection to each other) and analyze the performance of standard algorithms along these metrics in Section 7. As an example, if the follower runs a separate instantiation of ActiveArmElimination (Algorithm 7) on every arm $a\in\mathcal{A}$, this satisfies high-probability instantaneous regret $g(t,T,\mathcal{B})=O(\sqrt{|\mathcal{B}|\cdot\log T/t})$ and high-probability anytime regret $h(t,T,\mathcal{B})=O(\sqrt{|\mathcal{B}|\cdot t\cdot\log T})$ (Proposition 7.1).
2.4 Real-world examples
We describe two real-world examples which fit into our framework.
Example 2.1 (User-chatbot interaction).
Example 2.2 (User-recommender system interaction).
Consider interactions between a recommender system and a user, where recommender system gives a slate (or subset) of items $a\in\mathcal{A}$ to the user, and the user picks an action $b\in\mathcal{B}$ from the slate. When the user returns to the same content recommendation system (e.g. a Netflix/Hulu user with a profile) many times, this becomes a repeated game where both the recommendation system and user learn about their preferences (Hajiaghayi et al., 2023). Again, misalignment could occur from the engagement metric being misaligned with user welfare (Milli et al., 2021) or for unintentional reasons (e.g., misspecification due to discrete thumbs up/thumbs down user feedback, since true preferences are more nuanced).
Examples 2.1-2.2 motivate why our objective is to minimize regret for both the leader and the follower. First, we may inherently care about utility for the human, who could be either the leader (Example 2.1 in Section 2.4) or the follower (Example 2.2 in Section 2.4). Secondly, we may also care about utility for the algorithmic tools: for example, a recommendation system that fails to make money may go out of business, or in certain cases, the chatbot/recommender may better capture societal objective than certain humans.

### v2

@@ 1 Introduction
More broadly, our work takes a step towards assessing the utility of both learning agents in decentralized, misaligned environments. Our model and results capture the general setting where the player utilities are arbitrarily related, where players might not even agree upon which pairs of actions give similar or different rewards. This motivated us to design benchmarks which are tolerant to small errors in the other player. We hope that our benchmarks and algorithms serve as a starting point for assessing when two-agent learning systems in misaligned environments can ensure high utility for both agents.
@@ 1.1 Related Work
Several papers study decentralized online learning in SGs. Camara et al. (2020); Collina et al. (2023b) posit that the follower runs a no-counterfactual-internal-regret algorithm and design no-regret algorithms for the leader. However, they assume strong alignment between the players’ rewards: Camara et al. (2020) requires that a follower choosing an $\epsilon$-suboptimal action only results in an $O(\epsilon)$ utility loss for the leader. [^7] See Assumption 2 in Camara et al. (2020). Appendix D therein considers some relaxations, but they lead to $\Omega(T)$ worst-case regret. Collina et al. (2023b) partially relax this assumption, but still require the existence of stable actions for the leader. In contrast, we do not place any alignment conditions: in fact, misalignement is the driver of our linear regret result for the original Stackelberg benchmarks (Theorem 3.1). Other differences are that we focus on stochastic, rather than adversarial, rewards, and our benchmark is independent of the follower’s choice of learning algorithm. [^8] However, our regret bounds assume that the follower’s algorithm gracefully improves over time, see Section 2.4. Haghtalab et al. (2023) takes a different perspective and considers the follower running a calibrated algorithm. They design a leader algorithm which waits for the follower to partially converge, and show that the Stackelberg value is obtained in the limit as $T\rightarrow\infty$. In contrast, we focus on instance-independent regret bounds for a fixed time horizon $T$, which requires us to relax the benchmark. Other differences are we focus on stochastic, rather than deterministic, rewards, we assume the follower observes the leader’s action, and we consider the follower’s utility in addition to the leader’s utility.
@@
The literature on learning in SGs is vast and includes many other variations. Many works (e.g., (Letchford et al., 2009; Balcan et al., 2015; Zhao et al., 2023; Lauffer et al., 2023)) consider the leader performing (offline or) online learning and followers myopically best-responding. Other model variants studied include the leader strategizing against a follower who is running a no-regret learning algorithm (Braverman et al., 2018; Deng et al., 2019; Guruganesh et al., 2024; Brown et al., 2023; Lin and Chen, 2024), the leader and follower both running gradient-based algorithms (Fiez et al., 2019; Goktas et al., 2022), non-myopic followers who best-respond to a discounted utility over future time steps (Haghtalab et al., 2022; Hajiaghayi et al., 2023), repeated game formulations under complete information (Zuo and Tang, 2015; Collina et al., 2023a) the leader offering a menu of actions to the follower (Han et al., 2023), the (human) follower having cognitive biases in responding (Pita et al., 2009), and both players having side information (Harris et al., 2024). Other works have studied learning in structured SGs, including delegated choice (e.g., (Kleinberg and Kleinberg, 2018; Hajiaghayi et al., 2023)), strategic classification (e.g., (Dong et al., 2018; Chen et al., 2020; Zrnic et al., 2021; Ahmadi et al., 2021)), performative prediction (e.g., (Perdomo et al., 2020; Miller et al., 2021)), pricing under buyer and seller uncertainty (e.g., (Guo et al., 2023)), contract theory (e.g., (Zhu et al., 2023)), cake cutting (e.g., (Brânzei et al., 2024)), and aligned utilities (e.g., (Kao et al., 2022)).
@@
Our work also connects to a broader literature on interacting learners. This literature examines interactions between multiple bandit learners, studying aspects such as the convergence of systems of no-regret learners to coarse correlated and correlated equilibrium (e.g. (Daskalakis et al., 2011, 2021; Anagnostides et al., 2022)), multiple bandit learners competing for market share (e.g., (Aridor et al., 2020; Jagadeesan et al., 2023)), and multiple autobidding algorithms competing in an auction (e.g., (Borgs et al., 2007; Balseiro and Gur, 2019; Lucier et al., 2023)). Most closely related to this paper is corralling bandit algorithms (e.g., (Agarwal et al., 2017; Pacchiano et al., 2020)), where a “master algorithm” dynamically chooses among several “base algorithms”: our decentralized learning environment in the case of aligned player utilities is essentially an instance of corralling bandits, with the “base algorithms” corresponding to different leader actions. The interacting learner literature also examines human-algorithm collaboration studying aspects such as misalignment between engagement metrics and user welfare (e.g., (Ekstrand and Willemsen, 2016; Milli et al., 2021; Stray et al., 2021; Kleinberg et al., 2022)), impact of underspecification on human-AI misalignment (e.g., (Zhuang and Hadfield-Menell, 2020)), and “assistive” algorithmic tools (e.g. (Chan et al., 2019)). Most closely related to this paper is work on online learning in subset selection and conformal prediction, where goals often revolve around selecting a subset of items to present to a learning user (Straitouri and Rodriguez, 2023; Corvelo Benz and Rodriguez, 2024; Straitouri et al., 2023; Wang et al., 2022; Donahue et al., 2024; Agarwal and Brown, 2023; Yao et al., 2022), often with the goal of achieving complementarity (Bansal et al., 2021). The related area of human-AI interaction (see (Preece et al., 1994; Kim, 2015; MacKenzie, 2024; Lazar et al., 2017) for textbook treatments) studies similar questions, often from a more behavioral angle. More broadly, the interacting learner literature also studies applied domains including multi-agent reinforcement learning (see Zhang et al. (2019) for a survey) and federated learning (see Yang et al. (2019) for a survey).
@@ 2 Model and assumptions
In this section we describe our formal model. We first define an instance $\mathcal{I}=(\mathcal{A},\mathcal{B},v_{1},v_{2})$ in our setup, which captures the setup of the underlying static Stackelberg game. Let $\mathcal{A}$ be a finite action set for the leader (Player 1) and let $\mathcal{B}$ be a finite action set for the follower (Player 2). Let $v_{i}(a,b)\in[0,1]$ denote Player $i$’s value (i.e., mean reward) for the leader choosing $a$ and the follower choosing $b$. The Stackelberg equilibrium takes the following form. Let $b^{*}(a)$ be the best-response with respect to the follower’s rewards: [^9] In case of ties in follower utility, $b^{*}(a)$ is the best-response with lowest leader utility. $b^{*}(a)=\text{argmax}_{b\in\mathcal{B}}v_{2}(a,b)$. The Stackelberg equilibrium $(a^{*},b^{*})$ is defined to be the best action the leader can take, assuming that the follower will exactly best-respond:
@@
Note that for simplicity, we restrict both players to pure strategies. [^10] Other works (e.g., Bai et al. (2021)) similarly restricts both players to pure strategies.
In this paper, we move from the static Stackelberg Equilibrium environment to a repeated dynamic environment which we call a decentralized Stackelberg game (DSG). A DSG operates over $T$ time steps where each player selects actions using a multi-armed bandit algorithm. A DSG is sequential: at each time step $t$, the leader chooses an action $a_{t}\in\mathcal{A}$ and then the follower chooses an action $b_{t}\in\mathcal{B}$. A DSG is also decentralized: each player $i$ can observe their own stochastic rewards but not the stochastic rewards of the other player. We measure regret for each player $i$ by their cumulative reward across all time steps relative to a benchmark.
@@ 2.1 Interaction between players
In a DSG, the interaction between the leader and follower proceeds as follows. The leader chooses an algorithm $\texttt{ALG}_{1}$ mapping their history (formalized below) of observed actions and rewards to a distributions over actions $\mathcal{A}$, and the follower similarly chooses an algorithm $\texttt{ALG}_{2}$ mapping the leader’s action and the follower’s history to a distribution over actions $\mathcal{B}$. After the players select algorithms, the interaction between the leader and the follower proceeds as follows at each time step $t$:
@@
This interaction captures that the players are dynamic in their learning: in particular, this framework is sufficiently general to capture a wide range of learning strategies. However, we do not study the meta-game where agents strategically pick learning algorithms (e.g., see Kolumbus and Nisan (2022) for an example of a work studying the meta-game). We believe that our model captures many real-world environments such as user-chatbot interactions and recommender system-user interactions, where agents learn about their incurred rewards from past interactions even if they do not actively optimize the higher-order manner in which they learn. See Section 2.3 (Example 2.1 and Example 2.2) for more details of how these real-world examples are captured by our model.
@@
Having described how the players interact, we next discuss the players’ history, which further enforces decentralization. Each player $i$ can only observe their own reward $r_{i,t}(a_{t},b_{t})$ (and cannot observe the reward of the other player). In a strongly decentralized Stackelberg game (StrongDSG), the follower can observe the leader’s action $a_{t}$, but the leader cannot observe the follower’s action $b_{t}$, whereas in a weakly decentralized Stackelberg game (WeakDSG), the leader can additionally observe the follower’s action $b_{t}$. Notation for each player’s histories is presented in Section A.4.
Note that a strongly decentralized Stackelberg game (StrongDSG) restricts what information the leader has access to, and is thus more challenging than a weakly decentralized Stackelberg game (WeakDSG). One motivation for a strongly decentralized Stackelberg game is interpretability: the leader and follower may be taking actions in spaces that are not mutually understandable (e.g. a chatbot’s representation of human preferences may not be interpretable). Most of our positive results (i.e., algorithm constructions) focus on StrongDSGs, whereas our negative results apply to both StrongDSGs and WeakDSGs.
@@ 2.2 Measuring regret
Our goal is to obtain sublinear worst-case regret for both players: that is, we will assess $\max(R_{1}(T),R_{2}(T))$. We note that this challenging goal is a departure from previous work which has typically focused solely on sublinear regret for the leader (see Section 1.1). Our motivation for selecting this objective is that (a) a human could be either the leader or the follower, and (b) societal welfare may demand that we care about the utility of both the leader and the follower (discussed further in Section 2.3).
2.3 Real-world examples
We describe two real-world examples which fit into our framework.
Example 2.1 (User-chatbot interaction).
Example 2.2 (User-recommender system interaction).
Consider interactions between a recommender system and a user, where the recommender system gives a slate (or subset) of items $a\in\mathcal{A}$ to the user, and the user picks an action $b\in\mathcal{B}$ from the slate. When the user returns to the same content recommendation system (e.g. a Netflix/Hulu user with a profile) many times, this becomes a repeated game where both the recommendation system and user learn about their preferences (Hajiaghayi et al., 2023). Again, misalignment could occur from the engagement metric being misaligned with user welfare (Milli et al., 2021) or for unintentional reasons (e.g., misspecification due to discrete thumbs up/thumbs down user feedback, since true preferences are more nuanced).
Examples 2.1-2.2 motivate why our objective is to minimize regret for both the leader and the follower. First, we may inherently care about utility for the human, who could be either the leader (Example 2.1) or the follower (Example 2.2). Secondly, we may also care about utility for the algorithmic tools: for example, a recommendation system that fails to make money may go out of business, or in certain cases, the chatbot/recommender system may better capture societal objectives than certain humans.
@@ 2.4 Assumptions on the follower’s algorithm $\texttt{ALG}_{2}$
Finally, we present some technical assumptions on the follower’s algorithm. When we analyze regret in Sections 4-5, many of our constructions do not require that the follower run a particular algorithm, but instead allow the follower to run any algorithm that has sufficiently good performance along certain fine-grained performance metrics that capture the extent to which an algorithm’s performance gracefully improves over time.
@@
In a DSG, we will require similar properties to hold for the follower’s algorithm $\texttt{ALG}_{2}$, but we take account how the algorithm $\texttt{ALG}_{2}$ depends on the action $a_{t}$ which is selected by the leader’s algorithm $\texttt{ALG}_{1}$. Let $n_{t+1}(a)$ be the number of times that arm $a$ has been pulled prior to the $(t+1)$th time step. An algorithm $\texttt{ALG}_{2}$ satisfies a high-probability instantaneous regret bound of $g$ if for any $\texttt{ALG}_{1}$ chosen by the leader and any $\mathcal{I}=(\mathcal{A},\mathcal{B},v_{1},v_{2})$, it holds that:
$\mathbb{P}\left[\forall t\in[T],a\in\mathcal{A}\mid v_{2}(a_{t},b_{t})\geq\max_{b\in\mathcal{B}}v_{2}(a_{t},b)-g(n_{t+1}(a),T,\mathcal{B})\right]\geq 1-|\mathcal{A}|\cdot T^{-3}.$
@@
$\mathbb{P}\left[\forall t\in[T],a\in\mathcal{A}\mid\sum_{t^{\prime}\leq t\mid a_{t^{\prime}}=a}v_{2}(a,b_{t^{\prime}})\geq\left(\sum_{t^{\prime}\leq t\mid a_{t^{\prime}}=a}\max_{b\in\mathcal{B}}v_{2}(a,b)\right)-h(n_{t+1}(a),T,\mathcal{B})\right]\geq 1-|\mathcal{A}|\cdot T^{-3}.$
In Section 7, we discuss the relationship between these metrics, the performance of standard algorithms for the follower on these metrics, and algorithms for the leader for more general $g$ and $h$. As an example, if the follower runs a separate instantiation of ActiveArmElimination (Algorithm 7) on every arm $a\in\mathcal{A}$, this satisfies high-probability instantaneous regret $g(t,T,\mathcal{B})=O(\sqrt{|\mathcal{B}|\cdot\log T/t})$ and high-probability anytime regret $h(t,T,\mathcal{B})=O(\sqrt{|\mathcal{B}|\cdot t\cdot\log T})$ (Proposition 7.1).